Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report 2014-08-25

Observers: Leonardo Tartaglia, Nancy Elias-Rosa
Support: Stefano Benetti, Giacomo Terreran
TAT on duty: Maria Teresa Botticella, Andrea Pastorello

--------------------------------------------------------------

WEATHER CONDITIONS:
start: wind 9.3 m/s; temp. 10.0 C; humid. 15.9%; Clear.
middle: wind 13.5 m/s; temp. 11.8 C; humid. 7.0%; Clear.
end: wind 13.4 m/s; temp. 11.6 C; humid. 7.0%; Clear.


--------------------------------------------------------------

INSTRUMENTS: EFOSC2

CALIBRATIONS

Afternoon:
-- EFOSC2 bias, dome spec flats, arcs 

-- EFOSC2 skyflat UVi

Morning:
-- EFOSC2 skyflat BR

--------------------------------------------------------------

OBSERVATIONS

Target                                                    Type                  UT-Start         Filter/Grism           Slit (")         Seeing (")      Notes

LSQ14an                                                 Follow-up         23:25             UBVRi                     --              1.40 
CSS140424:133007-212728                  Follow-up         00:03             VRi                         --              0.78  
LSQ14eez                                               Follow-up         00:47             UBVRi+Gr#13         1.0             0.75  
LSQ14eer                                                Follow-up         03:01             Gr#11                    1.0             0.77             (1)   
SN2013hx                                               Follow-up        04:08              Ri                           --              0.97
LSQ14efd                                                Follow-up        04:51              UBVRi+Gr#13        1.0             1.07
VMa2                                                      Standard          07:36              Gr#11+Gr#13        1.0             0.85
OGLE-2014-SN-047                               Follow-up        07:50              UBVRi+Gr#13        1.0             1.04 
PSNJ01190082+1018191                       Classification    09:36              Gr#13                   1.0             1.00             (2)

(1) We rotate the slit to take both the SN and the center of the host galaxy in the same spectrum, as required by the PI. Unfortunately we have to stop the second spectrum 4 minutes earlier due to the increase of the wind speed. Now we are pointing limited. For this reason we did not take the photometry of this SN.
(2) We took a longer spectrum with better sky conditions than yesterday. 

Comments